OpenStudy (anonymous):

a diagonal length of a rectangular yard is 25 ft. the width of the yard is 10 feet what is the length of the yard rounded to the nearest tenth? choices are: 22. 9 feet 23.9 feet 25.9 feet or 26.9 feet expalain your answer

OpenStudy (anonymous):

You need to use the Pythagorean Theorem for this.

OpenStudy (anonymous):

\[hypotenuse = \sqrt{(x)^2 + (y)^2} = diagonal = \sqrt{(length)^2+(width)^2}\]

OpenStudy (anonymous):

? still a little unclear can you help with the process?

OpenStudy (anonymous):

See attached for steps. :)

OpenStudy (anonymous):

@lisamomhoffman